# Efficient Decoding Methods for Language Models on Encrypted Data

### v1

@@ 3.3 HE-Friendly Nucleus Sampling
In text generation, language models produce a probability distribution over a large vocabulary at each time step. The next token is selected from this distribution using a decoding strategy. The simplest strategy is greedy decoding, where the token with the highest probability ($\operatorname{argmax}$) is chosen. More commonly, however, high-quality text generation relies on stochastic decoding methods that introduce controlled randomness to improve fluency, diversity, and coherence. Two widely used stochastic decoding methods are $\operatorname{top-k}\operatorname{sampling}$ Fan et al. (2018) and $\operatorname{nucleus}$ (top-p) $\operatorname{sampling}$ Holtzman et al. (2019b). In $\operatorname{top-k}\operatorname{sampling}$, only the $k$ most probable tokens are considered, and one is sampled according to its probability. In $\operatorname{nucleus}\thinspace\operatorname{sampling}$, the model selects the smallest set of top-ranked tokens whose cumulative probability exceeds a threshold $p$, and samples from this adaptive set. These techniques prevent degeneration and repetition, making them essential in modern LLM-based systems.
Sampling over HE.
As established, our $\operatorname{CutMax}$ algorithm is polynomial and thus compatible with HE; we confirm its correctness empirically in Sect.˜4. Next, we propose efficient methods for sampling over HE, which build on top of our $\operatorname{CutMax}$. The first naive attempt is to use the inverse transform sampling method, which generates random samples from a given probability distribution using its cumulative distribution function (CDF): Given an encrypted probability density function (PDF) $X$ with CDF $F_{X}$, the method samples a plaintext value $U\sim\mathrm{Uniform}(0,1)$, and returns $x=F_{X}^{-1}(u)$, where $F_{X}^{-1}$ is the inverse CDF ($\mathrm{iCDF}$), and $x$ is a sample from $X$. This sampling method is inefficient as it requires at least one HE heavy comparison (Appendix˜B) to sample an element.
@@
Gumbel (1954) Let $U\sim{}\mathrm{Uniform}(0,1)$, then $G=-\log(-\log U)$ has a Gumbel distribution $G\sim{}\mathrm{Gumbel}(\mu=0,\beta=1)$.
Lemma 3.2 (Gumbel-Max Trick Maddison et al. (2014)).
@@
$\mathbb{P}(Y=i)=\frac{e^{X_{i}}}{\sum_{j}e^{X_{j}}}=Softmax(X_{i})\,.$
@@
While our Gumbel-based method allows efficient sampling over HE, we can do better. We aim to develop an algorithm inspired by Maddison et al. (2014) that will enable us to sample only from the top of the distribution, as in $\operatorname{nucleus}\thinspace\operatorname{sampling}$, so that words from the tail of the logits distribution could not be sampled, and potentially disrupt the coherence of the generated text. To this end, we aim to sample $G$ from a PDF that satisfy the condition $\mathbb{P}(G>p)=p$. Luckily, the $\mathrm{Beta}(\alpha,\beta)$ distribution has this property when $\beta=1$.
@@
Algorithm˜3 introduces a single-shot, nucleus-restricted sampling routine based on Lemma˜3.3. It starts by draw $G\sim{}\mathrm{Beta}(\alpha,\beta)$ (Line 4). Subsequently, we form the “cut-noise” logits at Line 5, so that only those tokens whose cumulative mass exceeds $p$ can potentially become the maximum. Finally, the algorithm executes a single invocation of $\operatorname{CutMax}$ at Line 6 to homomorphically recover the one-hot encrypted sample from the nucleus set. To summarize, we introduced a one-shot nucleus sampling algorithm that costs only one $\operatorname{CutMax}$ when evaluated over HE.
@@ 3.4 Differentiability for Sequence Optimization
This stands in contrast to non-differentiable ops like discrete $\operatorname{argmax}$ or $\operatorname{sampling}$, which require approximations like straight-through estimators (STE; Bengio et al. (2013)) during backpropagation—where gradients are passed ‘straight through’ the hard $\operatorname{argmax}$ as if it were a soft identity. STE introduces bias and can lead to optimization instabilities in sequence tasks (e.g., reinforcement learning from human feedback or sequence fine-tuning).
@@ 4 Empirical Evaluation
To evaluate our methods, we use the ccdv/arxiv-summarization Cohan et al. (2018) dataset of paper abstracts, sampling $N=100$ articles unless otherwise specified. We report results on realistic LLM logits from models including QWEN2.5-0.5B ($|\mathcal{V}|\approx 150\mathrm{K}$), Mistral-7B ($|\mathcal{V}|\approx 33\mathrm{K}$), and GPT-2 ($|\mathcal{V}|\approx 50\mathrm{K}$).
@@ 4.1 Plain-Text CutMax
We evaluate our $\operatorname{CutMax}$ in the unencrypted domain on the QWEN2.5-0.5B model (vocabulary size $|\mathcal{V}|\approx 150\mathrm{K}$), using the ccdv/arxiv-summarization Cohan et al. (2018) dataset of paper abstracts. We sampled $N=100$ articles and, for each, retrieved the next-token logits via one-step greedy decoding (greedy argmax on the plaintext logits), yielding 100 ground-truth tokens. For each prompt we then ran $\operatorname{CutMax}$ with $T\in\{2,3,4,5\}$ iterations, grid-searching over odd powers $p\in\{7,9,\dots,19\}$ and scaling factors $c\in\{3,5,\dots,39\}$ to find the smallest $(p,c)$ that still achieves perfect next-token recovery.
@@
Across all 400 $\operatorname{CutMax}$ invocations (100 prompts × 4 values of $T$), we achieve 100% next-token recovery. Moreover, after convergence (e.g., $T\geq 4$), the distribution is effectively one-hot: out of $\approx 150\mathrm{K}$ vocabulary items, only the true token carries non-negligible mass. These results confirm the theoretical analysis of Sect.˜3: $\operatorname{CutMax}$ drives a highly skewed distribution in just a handful of iterations, making it an excellent practical choice for homomorphic greedy decoding under CKKS-style encryption.
4.2 HE CutMax
We evaluated our FHE $\operatorname{CutMax}$ implementation under the CKKS scheme using HELayers (Aharoni et al., 2023). Experiments were conducted on LLM logits from two models: Qwen-2.5-0.5B and Mistral-7B, with vocabulary sizes of 151,936 and 32,768, respectively. For each model, we used 100 examples and set following hyperparameters: $c=[5,3,3,3],\quad p=[16,4,4,6],\quad T=4$, chosen to optimize the CKKS chain index and ensure numerical stability. We report the mean of the per-example maximum values, average latency, and slot prediction accuracy. As a baseline, we compare our method against the recently proposed HE-friendly $\operatorname{argmax}$ algorithm by Zhang et al. . Results are reported in Table˜2, demonstrating that our method is both accurate and efficient. From an accuracy perspective, our method does not alter the model’s predictions in any of the tested cases, achieving 100% accuracy. For efficiency benchmarking, our method significantly reduces the latency of $\operatorname{argmax}$ compared to Nexus Zhang et al. (2024a), by several orders of magnitude for comparable vocabulary sizes. Specifically, for a vocabulary size of approximately 30K, our method reduces latency from 79.56 seconds to 3.373 seconds. For a larger vocabulary of approximately 150K, our method reduces latency from 162.8 to 4.607. Note that we chose NEXUS as our baseline even though there are newer papers on LLM inference such as Zimerman et al. (2024a); Park et al. (2025); Zhang et al. (2025b); Moon et al. (2024b); Kei and Chow (2025), because these did not include an $\operatorname{argmax}$ benchmark or did not utilize $\operatorname{argmax}$ at all.
Table 2: Homomorphic $\operatorname{argmax}$ latency and accuracy. We benchmark $\operatorname{CutMax}$ on two realistic LLM logit tensors—Mistral-7B ($|\mathcal{V}|\!=\!32{,}768$) and Qwen-2.5-0.5B ($|\mathcal{V}|\!=\!151{,}936$)—using 100 prompts from the arXiv- summarisation set. Hyper-parameters are fixed to T=4 with p=[16,4,4,6] and c=[5,3,3,3] due to chain index optimization and numerical stability. We report the mean of the average of maximum values, the average latency, the accuracy of slot predictions, the arrays’ sizes, and the device used.
@@ 4.2 HE CutMax
Nexus Zhang et al. (2024a) Table 2: "We batched 32 inputs in total … Runtime is the amortized latency of each input.", originally reported (2.48,5.09)(2.48{,}5.09) when multiplied by 32 we get $(79.36,162.8)$, respectively.
@@ 4.3 Nucleus Sampling
Figure˜7 qualitatively demonstrates our method’s effectiveness, showing a sharply truncated distribution that confirms all probability mass remains within the top-$p$ nucleus. To verify that our one-shot nucleus sampler never selects tokens outside the intended top-$p$ set, we measured the violation rate as the fraction of draws falling outside the nucleus, over $S=100$ prompts, with $N=1000$ samples per prompt and $p=0.9$. We compare standard Gumbel-Max sampling against our Nucleus($\beta$–cut) method. Results are summarized in Figure˜4: Gumbel-Max violates the top-$p$ constraint in $9.3\%\pm 2.4\%$ of draws, whereas Nucleus ($\beta$-cut) has zero violations.
@@ 5 Conclusions
We solve a critical bottleneck in privacy-preserving AI: adapting LLM decoding for homomorphic encryption (HE), where non-polynomial operations like $\operatorname{argmax}$ and $\operatorname{sampling}$ are computationally prohibitive. We introduce CutMax, a novel, HE-friendly and theoretically grounded $\operatorname{argmax}$ algorithm that replaces slow, comparison-based methods with an efficient iterative polynomial process. Crucially, $\operatorname{CutMax}$ achieves 100% accuracy and executes in just a few seconds on large vocabularies ($|\mathcal{V}|\!\approx\!150\mathrm{K}$), making secure greedy decoding practical for LLM inference for the first time. Building on this, we present the first HE-compatible nucleus (top-p) sampling method, which leverages $\operatorname{CutMax}$ with the idea of noisy inverse transform sampling to provably restrict sampling to the desired token set. Together, these fully polynomial algorithms provide a complete and efficient framework for both greedy and stochastic decoding over encrypted data, addressing a major barrier to the deployment of secure LLMs in real-world applications.
@@ Appendix A Threat model
There are two commonly used threat-models for secure inference and generative AI for LLMs over HE: 1) using encrypted weights; or 2) using encrypted input samples (queries). When considering a 2-party scenario with HE, it is commonly assumed that one party is the (untrusted, semi-honest) server who holds the model (encrypted or not) and the other party is the data-owner who performs the query and would like to avoid revealing the query to the server. The decision of whether the server holds the model encrypted or not depends on whether a third-party the model-owner agrees to share the data with the server. Our proposed solution is orthogonal to the above decision, which eventually only affects latency. All communications are performed using TLS 1.3 and privacy attacks such as model extraction attack and membership inference attacks require extra mechanism such as differential privacy, which is also orthogonal to this work. For brevity we have decided to refer the reader to Adir et al. (2024)[Chapter 3] that further describes this security model.
@@ Appendix B Approximations for HE
where $\textsc{sign}(x)=1$ when $x>0$ and 1-1 otherwise. An approximation of sign is given, e.g., in Moon et al. (2024a). To make the results of sign suitable as an indicator function (as in Equation˜5), it is common to translate the sign range from $\{-1,1\}$ to $\{0,1\}$ using $\frac{1+\textsc{sign}(x)}{2}$. The $\max$ and sign functions are used by prior-art $\operatorname{argmax}$ implementations. In contrast, $\operatorname{CutMax}$ uses $\operatorname{Inv}$ and $\operatorname{InvSqr}$ approximations.
@@ B.1 Goldschmidt
The Goldschmidt inverse algorithm Goldschmidt (1964) is an iterative process for computing the function $f(x)=\frac{1}{x}$. Cheon et al. (2019) showed that this algorithm can be used to approximate inputs in the range $[0,2]$ effectively using the equation
@@
Specifically, they proved that for $(x\in\left[\frac{1}{2^{n}},1\right])$, it requires $(d=\Theta(\log\alpha+n))$ iterations to converge, with an error bound of $(2^{\alpha})$. Panda (2021) showed an efficient implementation for the Goldschmidt inverse square root ($\operatorname{InvSqr}$) algorithm over HE. Subsequently, Moon et al. (2024a) reported improved measurements of the inverse and $\operatorname{InvSqr}$ implementations under CKKS. Table˜3 repeats some benchmarks for the sign and $\operatorname{InvSqr}$ functions at different approximation levels ($\alpha$). One immediate conclusion is that the $\operatorname{InvSqr}$ implementation is $(1.5-2.1\times)$ faster compared to the sign implementation for the same level of errors. One reason is that $\operatorname{InvSqr}$ has a lower multiplication depth (defined below) and thus requires fewer bootstraps.
@@
Table 3: A sample benchmarks of Moon et al. (2024a) conducted on AMD EPYC 7502-32 CPU, using Microsoft Seal, poly degree of $2^{17}$, and scale $\Delta=2^{40}$.

### v2

@@ 3.3 HE-Friendly Nucleus Sampling
In text generation, language models produce a probability distribution over a large vocabulary at each time step. The next token is selected from this distribution using a decoding strategy. The simplest strategy is greedy decoding, where the token with the highest probability ($\operatorname{argmax}$) is chosen. More commonly, however, high-quality text generation relies on stochastic decoding methods that introduce controlled randomness to improve fluency, diversity, and coherence. Two widely used stochastic decoding methods are $\operatorname{top-k}\operatorname{sampling}$ Fan_2018 and $\operatorname{nucleus}$ $(\operatorname{top-p})$ $\operatorname{sampling}$ Holtzman_2020. In $\operatorname{top-k}\operatorname{sampling}$, only the $k$ most probable tokens are considered, and one is sampled according to its probability. In $\operatorname{nucleus}\thinspace\operatorname{sampling}$, the model selects the smallest set of top-ranked tokens whose cumulative probability exceeds a threshold $p$, and samples from this adaptive set. These techniques prevent degeneration and repetition, making them essential in modern LLM-based systems.
Sampling under HE.
As established, our $\operatorname{CutMax}$ algorithm is polynomial and thus compatible with HE; we confirm its correctness empirically in Sect.˜4. Whereas $\operatorname{CutMax}$ handles deterministic decoding, we next extend it to stochastic sampling. The first naive attempt is to use the inverse transform sampling method, which generates random samples from a given probability distribution using its cumulative distribution function (CDF): Given an encrypted probability density function (PDF) $X$ with CDF $F_{X}$, the method samples a plaintext value $U\sim\mathrm{Uniform}(0,1)$, and returns $x=F_{X}^{-1}(u)$, where $F_{X}^{-1}$ is the inverse CDF ($\mathrm{iCDF}$), and $x$ is a sample from $X$. This sampling method is inefficient as it requires at least one HE heavy comparison (Appendix˜B) to sample an element.
@@
gumbel1954statistical Let $U\sim{}\mathrm{Uniform}(0,1)$, then $G=-\log(-\log U)$ has a Gumbel distribution $G\sim{}\mathrm{Gumbel}(\mu=0,\beta=1)$.
Lemma 3.2 (Gumbel-Max Trick maddison2014).
@@
$\mathbb{P}(Y=i)=\frac{e^{X_{i}}}{\sum_{j}e^{X_{j}}}=\operatorname{Softmax}(X_{i})\,.$
@@
While our Gumbel-based method allows efficient sampling over HE, we can do better. We aim to develop an algorithm inspired by maddison2014 that will enable us to sample only from the top of the distribution, as in $\operatorname{nucleus}\thinspace\operatorname{sampling}$, so that words from the tail of the logits distribution could not be sampled, and potentially disrupt the coherence of the generated text. To this end, we aim to sample $G$ from a PDF that satisfy the condition $\mathbb{P}(G>p)=p$. Luckily, the $\mathrm{Beta}(\alpha,\beta)$ distribution has this property when $\beta=1$.
@@
Algorithm˜3 introduces a single‐shot, nucleus-restricted sampling routine based on Lemma˜3.3. It starts by draw $G\sim{}\mathrm{Beta}(\alpha,\beta)$ (Line 4). Subsequently, we form the “cut-noise” logits at Line 5, so that only those tokens whose cumulative mass exceeds $p$ can potentially become the maximum. Finally, the algorithm executes a single invocation of $\operatorname{CutMax}$ at Line 6 to homomorphically recover the one-hot encrypted sample from the nucleus set. To summarize, we introduced a one-shot nucleus sampling algorithm whose polynomial formulation requires only a single $\operatorname{CutMax}$ evaluation under HE, while maintaining exact distributional behavior and full HE compatibility.
@@ 3.4 Differentiability for Sequence Optimization
This stands in contrast to non-differentiable ops like discrete $\operatorname{argmax}$ or $\operatorname{sampling}$, which require approximations like straight-through estimators (STE; Bengio2013EstimatingOP) during backpropagation—where gradients are passed ‘straight through’ the hard $\operatorname{argmax}$ as if it were a soft identity. STE introduces bias and can lead to optimization instabilities in sequence tasks (e.g., reinforcement learning from human feedback or sequence fine-tuning).
@@ 4 Empirical Evaluation
To evaluate our methods, we use the ccdv/arxiv-summarization cohan-etal-2018-discourse dataset of paper abstracts, sampling $N=100$ articles unless otherwise specified. We report results on realistic LLM logits from models including QWEN2.5-0.5B ($|\mathcal{V}|\approx 150\mathrm{K}$), Mistral-7B ($|\mathcal{V}|\approx 33\mathrm{K}$), and GPT-2 ($|\mathcal{V}|\approx 50\mathrm{K}$).
@@ 4.1 Plain-Text CutMax
We evaluate our $\operatorname{CutMax}$ in the unencrypted domain on the QWEN2.5-0.5B model (vocabulary size $|\mathcal{V}|\approx 150\mathrm{K}$), using the ccdv/arxiv-summarization cohan-etal-2018-discourse dataset of paper abstracts. We sampled $N=100$ articles and, for each, retrieved the next‐token logits via one‐step greedy decoding (greedy argmax on the plaintext logits), yielding 100 ground‐truth tokens. For each prompt we then ran $\operatorname{CutMax}$ with $T\in\{2,3,4,5\}$ iterations, grid‐searching over odd powers $p\in\{7,9,\dots,19\}$ and scaling factors $c\in\{3,5,\dots,39\}$ to find the smallest $(p,c)$ that still achieves perfect next‐token recovery.
@@
Across all 400 $\operatorname{CutMax}$ invocations (100 prompts × 4 values of $T$), we achieve 100% next‐token recovery. Moreover, after convergence (e.g., $T\geq 4$), the distribution is effectively one‐hot: out of $\approx 150\mathrm{K}$ vocabulary items, only the true token carries non-negligible mass. These results confirm the theoretical analysis of Sect.˜3: $\operatorname{CutMax}$ drives a highly skewed distribution in just a handful of iterations, making it an excellent practical choice for homomorphic greedy decoding under CKKS‐style encryption.
Table 2: Homomorphic $\operatorname{argmax}$ latency and accuracy. We benchmark $\operatorname{CutMax}$ on two realistic LLM logit tensors—Mistral‑7B ($|\mathcal{V}|\!=\!32{,}768$) and Qwen‑2.5‑0.5B ($|\mathcal{V}|\!=\!151{,}936$)—using 100 prompts from the arXiv‑ summarisation set. Hyper‑parameters are fixed to T=4 with p=[16,4,4,6] and c=[5,3,3,3] due to chain index optimization and numerical stability. We report the mean of the average of maximum values, the average latency, the accuracy of slot predictions, the arrays’ sizes, and the device used.
@@
Nexus transformerArgmax Table 2: "We batched 32 inputs in total … Runtime is the amortized latency of each input.", originally reported (2.48,5.09)(2.48{,}5.09) when multiplied by 32 we get $(79.36,162.8)$, respectively.
4.2 HE CutMax
We evaluated our FHE $\operatorname{CutMax}$ implementation under the CKKS scheme using HELayers (helayers). Experiments were conducted on LLM logits from two models: Qwen-2.5-0.5B and Mistral-7B, with vocabulary sizes of 151,936 and 32,768, respectively. For each model, we used 100 examples and set following hyperparameters: $c=[5,3,3,3],\quad p=[16,4,4,6],\quad T=4$, chosen to optimize the CKKS chain index and ensure numerical stability. We report the mean of the per-example maximum values, average latency, and slot prediction accuracy. As a baseline, we compare our method against the recently proposed HE-friendly $\operatorname{argmax}$ algorithm by zhang2024secure. Results are reported in Table˜2, demonstrating that our method is both accurate and efficient. From an accuracy perspective, our method does not alter the model’s predictions in any of the tested cases, achieving 100% accuracy. For efficiency benchmarking, our method significantly reduces the latency of $\operatorname{argmax}$ compared to Nexus transformerArgmax, by several orders of magnitude for comparable vocabulary sizes. Specifically, for a vocabulary size of approximately 30K, our method reduces latency from 79.56 seconds to 3.373 seconds. For a larger vocabulary of approximately 150K, our method reduces latency from 162.8 to 4.607. Note that we chose Nexus as our baseline even though there are newer papers on LLM inference such as zimerman2024power; powerformer; moai; thor; kei2025shaft, because these did not include an $\operatorname{argmax}$ benchmark or did not utilize $\operatorname{argmax}$ at all.
@@ 4.3 Nucleus Sampling
Figure˜5 qualitatively demonstrates our method’s effectiveness, showing a sharply truncated distribution that confirms all probability mass remains within the top-$p$ nucleus. To verify that our one‐shot nucleus sampler never selects tokens outside the intended top-$p$ set, we measured the violation rate as the fraction of draws falling outside the nucleus, over $S=100$ prompts, with $N=1000$ samples per prompt and $p=0.9$. We compare standard Gumbel-Max sampling against our Nucleus($\beta$–cut) method. Results are summarized in Figure˜4: Gumbel-Max violates the top-$p$ constraint in $9.3\%\pm 2.4\%$ of draws, whereas Nucleus ($\beta$‐cut) has zero violations.
@@ 5 Conclusions
We solve a critical bottleneck in privacy-preserving AI: adapting LLM decoding for homomorphic encryption, where non-polynomial operations like $\operatorname{argmax}$ and $\operatorname{sampling}$ are computationally prohibitive. We introduce CutMax, a novel, HE-friendly and theoretically grounded $\operatorname{argmax}$ algorithm that replaces slow, comparison-based methods with an efficient iterative polynomial process. Crucially, $\operatorname{CutMax}$ achieves 100% accuracy and executes in just a few seconds on large vocabularies ($|\mathcal{V}|\!\approx\!150\mathrm{K}$), making secure greedy decoding practical for LLM inference for the first time. Building on this, we present the first HE-compatible nucleus (top-p) sampling method, which leverages $\operatorname{CutMax}$ with the idea of noisy inverse transform sampling to provably restrict sampling to the desired token set. Together, these fully polynomial algorithms provide a complete and efficient framework for both greedy and stochastic decoding over encrypted data, addressing a major barrier to the deployment of secure LLMs in real-world applications.
@@
Figure 5: Comparison of Sampling Methods on GPT-2 Last-Token Decoding: The top strip shows the complete 50,257‑dimensional logit vector (only its outline is visible at this scale), and the second strip zooms in on the five largest logits. The three lower strips plot, for 1,000 repeated draws, histograms of the 30 most frequently chosen token indices produced by four different samplers: standard $\operatorname{Softmax}$ at temperature 1, the classical Gumbel‑Max trick, inverse‑transform sampling on the normalized exponentials (“Exp‑inverse”), and our one‑shot Beta‑cut nucleus sampler, which perturbs each logit with i.i.d. $G_{i}\!\sim\!\mathrm{Beta}(\alpha,1)$ noise using $\alpha=0.486$ computed from ˜3.4 for $(p,q)=(0.9,0.95)$. Whereas the first three methods still allocate non‑negligible mass to tail tokens, the Beta‑cut histogram is sharply truncated, confirming that all probability remains inside the top‑$p$ nucleus exactly as required by the HE‑friendly design of Sect.˜$3.$
@@ Appendix A Threat model
There are two commonly used threat-models for secure inference and generative AI for LLMs over HE: 1) using encrypted weights; or 2) using encrypted input samples (queries). When considering a 2-party scenario with HE, it is commonly assumed that one party is the (untrusted, semi-honest) server who holds the model (encrypted or not) and the other party is the data-owner who performs the query and would like to avoid revealing the query to the server. The decision of whether the server holds the model encrypted or not depends on whether a third-party the model-owner agrees to share the data with the server. Our proposed solution is orthogonal to the above decision, which eventually only affects latency. All communications are performed using TLS 1.3 and privacy attacks such as model extraction attack and membership inference attacks require an extra mechanism such as differential privacy, which is also orthogonal to this work. For brevity, we have decided to refer the reader to he4ds[Chapter 3] that further describes this security model.
@@ Appendix B Approximations for HE
where $\textsc{sign}(x)=1$ when $x>0$ and 1-1 otherwise. An approximation of sign is given, e.g., in asor. To make the results of sign suitable as an indicator function (as in Equation˜5), it is common to translate the sign range from $\{-1,1\}$ to $\{0,1\}$ using $\frac{1+\textsc{sign}(x)}{2}$. The $\max$ and sign functions are used by prior-art $\operatorname{argmax}$ implementations. In contrast, $\operatorname{CutMax}$ uses $\operatorname{Inv}$ and $\operatorname{InvSqr}$ approximations.
@@ B.1 Goldschmidt
The Goldschmidt inverse algorithm goldschmidt is an iterative process for computing the function $f(x)=\frac{1}{x}$. goldapprox showed that this algorithm can be used to approximate inputs in the range $[0,2]$ effectively using the equation
@@
Specifically, they proved that for $(x\in\left[\frac{1}{2^{n}},1\right])$, it requires $(d=\Theta(\log\alpha+n))$ iterations to converge, with an error bound of $(2^{\alpha})$. panda showed an efficient implementation for the Goldschmidt inverse square root ($\operatorname{InvSqr}$) algorithm over HE. Subsequently, asor reported improved measurements of the inverse and $\operatorname{InvSqr}$ implementations under CKKS. Table˜3 repeats some benchmarks for the sign and $\operatorname{InvSqr}$ functions at different approximation levels ($\alpha$). One immediate conclusion is that the $\operatorname{InvSqr}$ implementation is $(1.5-2.1\times)$ faster compared to the sign implementation for the same level of errors. One reason is that $\operatorname{InvSqr}$ has a lower multiplication depth (defined below) and thus requires fewer bootstraps.
@@
Table 3: A sample benchmarks of asor conducted on AMD EPYC 7502-32 CPU, using Microsoft Seal, poly degree of $2^{17}$, and scale $\Delta=2^{40}$.